# Retrieval-Augmented Generation: Is Dense Passage Retrieval Retrieving?

### v1

@@ 1 Introduction
In just a few years, Large Language Models (LLMs) have emerged from the research labs to become a tool utilized daily by hundreds of millions of people and integrated into a wide variety of businesses. Despite their popularity, these models have been critiqued for frequently hallucinating, confidently outputting incorrect information Bang et al. (2023). Such inaccuracies not only mislead people but also erode trust in LLMs. Trust in these systems is crucial to their success and rate of adoption.
@@
Retrieval methods can be broadly categorized into two types: sparse and dense Zhao et al. (2024). Sparse methods encode queries and passages into sparse vectors, usually based on terms that appear in the queries and passages Robertson and Zaragoza (2009); Sparck Jones (1972). Dense methods employ language models to encode the semantic information in queries and passages into dense vectors Karpukhin et al. (2020); Huang et al. (2013). Dense methods often share two common properties: (a) the joint training of two or more encoding models – one for embedding a query and the other for embedding a knowledge base, and (b) contrastive training. These commonalities were introduced in the dense passage retrieval method, inspiring many subsequent methods in the literature.
In this paper, we analyze the original DPR method using the BERT-base backbone. We begin by analyzing the effects of training and using different layers for retrieval (Section 2). Then we continue by probing the model to determine if the features of pre-trained BERT are as discriminative as DPR-BERT in matching a query to the correct hard-negative passage (Section 3). Next, using techniques from the pruning literature, we compare the relative strength and number of activations of the feedforward layers throughout the original pre-trained and DPR-trained models (Section 4). Finally, we add and remove knowledge from the network to investigate how knowledge interacts with DPR training (Section 5). Through these experiments, we analyze DPR from multiple perspectives to understand what is changing in the backbone model during the training process.
2 Early Passage Match Capability
Figure 1: Per layer accuracy of the baseline DPR model and percentage of incorrectly matched output that was correctly matched at layer N.
Multilayer transformers are typically used for DPR. These models perform layer-to-layer updates to refine and promote concepts within the embedding space Geva et al. (2022). Such iterative refinement can reach a point of saturation, beyond which additional updates in later layers contribute marginally to model performance. This saturation suggests that there may be enough information early in the model to accurately match a query to a passage.
In this section, we analyze model update saturation to identify: (a) which layers contribute most significantly to performance improvements, and (b) aside from the final layer, which layers are crucial for training. Our analysis employs the training methodology outlined in the earlier DPR paper Karpukhin et al. (2020). The BERT model was fine-tuned in this style on the Natural Questions (NQ) dataset Kwiatkowski et al. (2019). To assess the DPR model’s accuracy at layer N, we conducted a maximum inner product search using the hidden state of layer N.
Due to the number of different training configurations and layers that were tested, it was not possible to re-index the 21M passage Wikipedia knowledge base that is commonly used with NQ. Therefore, we focused on assessing in-batch match accuracy. Each batch comprised 64 queries and 832 passages. The performance metrics reported in this section reflect the accuracy rate, namely the proportion of queries that were successfully matched to a corresponding passage within the batch.
Figure 2: Accuracy of DPR model when only layer N and layer 12 are trained with the full NQ dataset and only 30% of the NQ dataset
Figure 3: Percent of incorrect matches that intermediary network layers got correct.
Task
Model
Layer 0
Layer 1
Layer 2
Layer 3
Layer 4
Layer 5
Layer 6
Layer 7
Layer 8
Layer 9
Layer 10
Layer 11
Layer 12
2-Passage Probing
Pre-trained BERT – Untrained Probe
0.50
0.50
0.51
0.48
0.50
0.52
0.51
0.51
0.50
0.49
0.50
0.54
0.50
Pre-trained BERT
0.51
0.69
0.74
0.74
0.77
0.79
0.81
0.81
0.81
0.82
0.83
0.84
0.84
DPR-BERT Query Model
0.51
0.68
0.74
0.77
0.79
0.80
0.81
0.83
0.82
0.83
0.83
0.82
0.82
DPR-BERT Context Model
0.51
0.68
0.74
0.77
0.79
0.80
0.81
0.83
0.82
0.83
0.83
0.82
0.82
3-Passage Probing
Pre-trained BERT
0.34
0.53
0.59
0.59
0.65
0.64
0.67
0.67
0.68
0.69
0.69
0.73
0.73
DPR-BERT
0.34
0.54
0.60
0.63
0.66
0.66
0.66
0.70
0.71
0.69
0.73
0.72
0.71
4-Passage Probing
Pre-trained BERT
0.26
0.43
0.47
0.49
0.53
0.57
0.61
0.60
0.56
0.62
0.64
0.66
0.66
DPR-BERT
0.26
0.46
0.51
0.54
0.57
0.58
0.60
0.63
0.64
0.63
0.65
0.63
0.63
5-Passage Probing
Pre-trained BERT
0.21
0.35
0.42
0.43
0.43
0.50
0.53
0.53
0.54
0.56
0.57
0.60
0.61
DPR-BERT
0.21
0.36
0.42
0.48
0.49
0.51
0.54
0.56
0.58
0.58
0.60
0.56
0.56
@@ 2 Early Passage Match Capability
Figure 1 illustrates that peak performance is achieved at the network’s final layer. Notably, the most substantial performance increases occur between layers 3 and 4, 9 and 10, and 10 and 11. Moreover, Figure 1 suggests that in some cases, layers 4, 10, and 11 had the highest occurrence where the intermediary hidden layers correctly identified the relevant passage but a later layer subsequently revised the prediction to an incorrect match.
To further investigate the relationship between individual model layers and DPR performance, a series of experiments were conducted where the DPR model was trained two layers at a time. In these experiments, all layers were frozen except for two: layer N, which progressed from the first to the last layer across subsequent experiments, and the final layer, included due to its superior performance. Training followed the protocol established in the DPR paper, with layers N and 12 being trained independently—resulting in two separate loss calculations per batch, one for the interaction between the hidden states at layer N in the query and passage models and one loss for the embeddings outputted at layer 12 in the query and passage models.
Trained over the Natural Questions (NQ) dataset Kwiatkowski et al. (2019), the experiment demonstrated optimal performance when layers 5, 7, or 8 were concurrently trained with layer 12, as shown in Figure 2. However, since the performance gain was only slight, the experiment was repeated under a data-restricted regime, utilizing only $30\%$ of the NQ dataset. In this constrained data scenario, the disparity in accuracy when training various layers became more pronounced. Notably, training layer 7 provided the greatest accuracy boost. Moreover, training any layer within the middle segment of the network (layers 5-8) surpassed the performance of layers closer to the beginning or end of the network.
Figures 2 and 3 demonstrate that training just one middle layer in conjunction with layer 12 not only yields the highest performance but also leads to a middling frequency of instances where a previously correct match is incorrectly altered later in the model. According to prior research, it is in these middle layers that the processing of syntactic and semantic features co-occur Geva et al. (2021). The syntactic features are more akin to sparse retrieval methods whereas the semantic features represent the theme of DPR-style retrieval methods. This suggests that learning the interplay between syntactic and semantic features is likely an important part of what DPR-style training teaches the model.
@@ 3 Knowledge Consistency between untrained and trained model
Table 1 shows the result of this experiment. The performance disparity between probes for pre-trained BERT and DPR-trained BERT is relatively minor in the two-passage scenario ($1.8\%$) and interestingly, it is the pre-trained BERT that exhibits a slight advantage. As the number of passages increases, the performance gap widens to approximately $6\%$, and overall probe efficacy declines. These findings suggest that the inherent capabilities to discern relevant from irrelevant passages are already present in pre-trained BERT, and DPR-style training does not substantially enhance these discriminative features. Moreover, echoing the previous section’s findings, the hidden features in the network’s middle layers perform nearly as well as those from the final layers.
@@ 4 Knowledge Decentralization in DPR-Trained Models
The Riemann approximation was used due to the intractability of calculating a continuous integral. Following Dai et al. (2022), a threshold of $0.1*\max(\textrm{Attr})$ was applied to identify a coarse set of knowledge neurons. Appendix A.2 demonstrates that our observations are consistent across a spectrum of thresholds. In contrast to Dai et al. (2022), the coarse set of knowledge neurons was not refined to a fine set of knowledge neurons, as our interest is on the broader activation patterns. When the model operates, it activates both "true-positive" and "false-positive" knowledge neurons indiscriminately according to their attribution scores. The primary interest lies in how DPR training influences these activation patterns, rather than the role of specific neurons.
@@
Figure 4 illustrates the impact of DPR training on BERT’s neuron activations, charting the attribution score of every neuron across both the intermediate and output linear layers within each transformer block for the query model. Appendix A.1 shows that the observations made in this section also hold for the context model. DPR-trained BERT has more activated neurons in the intermediate layer of each block. The output layer, on the other hand, maintains a consistent number of activations at each transformer block compared to pre-trained BERT, and in the earlier layers DPR-trained BERT activates fewer neurons in the output layers. Previous studies have conceptualized intermediate layers as "keys" and the output layer as the "value" Geva et al. (2021). This suggests that DPR training expands the set of "keys" available to access a given volume of semantic knowledge while decreasing the accessible volume of syntactic knowledge, embodying a decentralization strategy for semantic knowledge. Rather than relying on a single, highly precise key to unlock some knowledge, DPR allows for the use of multiple, somewhat less precise keys. This underscores DPR training’s primary goal: to modify the model’s method of knowledge access without altering the stored knowledge itself. These multiple pathways enable morphologically distinct but semantically related text to trigger the same knowledge or collections of facts, thus making retrieval possible.
@@ 5.2 Knowledge Removal
Transformer-Patch
0.16
689
0.87
183
MalMen
0.11
721
0.81
261
Mend
0.11
722
1.00
252
Table 3: This table presents the outcomes of the knowledge removal experiments. The "Probing Removed" column is the percentage of the total facts that were successfully added to BERT. The "DPR Removed" column is the percentage of those facts that were detected after DPR training.
@@ 6 Related Works
Dense passage retrieval (DPR) addresses the challenge of matching a query with the most relevant passages from a knowledge base Karpukhin et al. (2020). This approach employs dual encoders—one encoder for the passages and another for the query—and utilizes a distance metric, such as the inner product, to identify the passages closest to the query. Inspired by Siamese networks Bromley et al. (1993), DPR represents the first fully neural architecture to outperform the BM25 algorithm Robertson and Zaragoza (2009). Since then, there have been quite a few improvements in how to train DPR-style models. Methods like RocketQA improve DPR by employing cross-batch negatives and training the network on more difficult hard negatives Qu et al. (2021). Dragon focuses on novel data augmentation and supervision strategies Lin et al. (2023). Contriever also employs a greater number of hard-negatives and data-augmentation methods in addition to pre-training the model on the inverse cloze task Izacard et al. (2022). MVR generates multiple views for each document to allow for multiple diverse representations of each of them Zhang et al. (2022). ColBERT employs token embeddings for more fine-grained matching Khattab and Zaharia (2020). REALM leverages feedback from the reader component to jointly train the retriever with the reader Guu et al. (2020). Other methods distill knowledge from the reader to the retriever Izacard and Grave (2020); Reichman and Heck (2023). Additionally, efforts in query augmentation or generation aim to better synchronize the query with the document encoder Ma et al. (2023); Wang et al. (2023); Shao et al. (2023); Gao et al. (2023). Despite these different enhancements, each method builds upon the DPR framework discussed in this paper.
@@ 7 Conclusion
To reveal possible avenues for improving retrieval augmented generation (RAG) systems, this paper set out to study the purpose served by DPR-style fine-tuning and how DPR-trained BERT operates. We found in Section 2 that middle layers of the LLM impact the performance the most. These layers are where the model is processing a mix of syntactic and semantic features according to Geva et al. (2021). Through linear probing in Section 3, alongside experiments where we added and removed knowledge from pre-trained BERT in Section 5, we determined that BERT does not appear to acquire new information through DPR fine-tuning. Instead, we observed that the efficacy of retrieval hinges on the activation of shared facts/memories between the BERT models used to encode the query and the context passages. This mechanism implies that incorrect retrieval could occur if a query or context passage inadvertently activates irrelevant or incorrect memories in BERT. Moreover, the absence of necessary facts or webs of knowledge within the model hampers its ability to retrieve information.
@@
In the most fundamental sense, Dense Passage Retrieval achieves its namesake function—it retrieves, locating and returning relevant context to the user given a query. Yet, as our evidence suggests, DPR models appear constrained to retrieving information based on the knowledge that preexists within their parameters, either innately or through augmentation. This operational boundary delineates a significant caveat: facts must already be encoded within the model for useful context to be accessible by retrieval. Absent these facts or their associative networks, retrieval seems to falter. Thus, if retrieval is understood as the capacity to recall or recognize knowledge already familiar to the model, then indeed, DPR models fulfill this criterion. However, if we extend our definition of retrieval to also encompass the ability to navigate and elucidate concepts previously unknown or unencountered by the model—a capacity akin to how humans research and retrieve information—our findings imply that DPR models fall short of this mark.

### v2

@@ 1 Introduction
In just a few years, Large Language Models (LLMs) have emerged from research labs to become a tool utilized daily by hundreds of millions of people and integrated into a wide variety of businesses. Despite their popularity, these models have been critiqued for frequently hallucinating, confidently outputting incorrect information Bang et al. (2023). Such inaccuracies not only mislead people but also erode trust in LLMs. Trust in these systems is crucial to their success and rate of adoption.
@@
Task Model Layer 0 Layer 1 Layer 2 Layer 3 Layer 4 Layer 5 Layer 6 Layer 7 Layer 8 Layer 9 Layer 10 Layer 11 Layer 12
2-Passage Probing Pre-trained BERT – Untrained Probe 0.50 0.50 0.51 0.48 0.50 0.52 0.51 0.51 0.50 0.49 0.50 0.54 0.50
Pre-trained BERT 0.51 0.69 0.74 0.74 0.77 0.79 0.81 0.81 0.81 0.82 0.83 0.84 0.84
DPR-BERT Query Model 0.51 0.68 0.74 0.77 0.79 0.80 0.81 0.83 0.82 0.83 0.83 0.82 0.82
DPR-BERT Context Model 0.51 0.68 0.74 0.77 0.79 0.80 0.81 0.83 0.82 0.83 0.83 0.82 0.82
3-Passage Probing Pre-trained BERT 0.34 0.53 0.59 0.59 0.65 0.64 0.67 0.67 0.68 0.69 0.69 0.73 0.73
DPR-BERT 0.34 0.54 0.60 0.63 0.66 0.66 0.66 0.70 0.71 0.69 0.73 0.72 0.71
4-Passage Probing Pre-trained BERT 0.26 0.43 0.47 0.49 0.53 0.57 0.61 0.60 0.56 0.62 0.64 0.66 0.66
DPR-BERT 0.26 0.46 0.51 0.54 0.57 0.58 0.60 0.63 0.64 0.63 0.65 0.63 0.63
5-Passage Probing Pre-trained BERT 0.21 0.35 0.42 0.43 0.43 0.50 0.53 0.53 0.54 0.56 0.57 0.60 0.61
DPR-BERT 0.21 0.36 0.42 0.48 0.49 0.51 0.54 0.56 0.58 0.58 0.60 0.56 0.56
@@
Retrieval methods can be broadly categorized into two types: sparse and dense Zhao et al. (2024). Sparse methods encode queries and passages into sparse vectors, usually based on terms that appear in the queries and passages Robertson and Zaragoza (2009); Sparck Jones (1972). Dense methods employ language models to encode the semantic information in queries and passages into dense vectors Karpukhin et al. (2020); Huang et al. (2013). Dense methods often share two common properties: (a) the joint training of two or more encoding models – one for embedding a query and the other for embedding a knowledge base, and (b) contrastive training. These commonalities were introduced in the DPR method, inspiring many subsequent methods in the literature.
In this paper, we analyze the original DPR method using the BERT-base backbone. We begin by probing the model to determine if the features of pre-trained BERT are as discriminative as DPR-BERT in matching a query to the correct passage amongst hard-negative passages (Section 2). Next, using techniques from the pruning literature, we compare the relative strength and number of activations of the feedforward layers throughout the original pre-trained and DPR-trained models (Section 3). Finally, we add and remove knowledge from the network to investigate how knowledge interacts with DPR training (Section 4). Through these experiments, we analyze DPR from multiple perspectives to understand what is changing in the backbone model during the training process.
@@ 2 Knowledge Consistency Between Untrained and Trained Model
The difference between a true positive passage and a hard-negative passage is usually the presence of 1-2 key distinct facts in the passage. The ability to discriminate between 2-5 of these passages indicates that the model likely has the awareness of which facts are relevant to the query. This awareness is likely driven by the model’s knowledge of the subject (as discussed in later sections of this paper). Rather than testing overall retrieval ability, this experiment aims to find how aware/knowledgeable pre-trained BERT’s features are compared to DPR-trained BERT when the difference of knowing or not knowing 1-2 facts can impact the final matching prediction.
Table 1 shows the result of this experiment. The performance disparity between probes for pre-trained BERT and DPR-trained BERT is relatively minor in the two-passage scenario ($1.8\%$) and interestingly, it is the pre-trained BERT that exhibits a slight advantage. As the number of passages increases, the performance gap widens to approximately $6\%$, and overall probe efficacy declines. These findings suggest that the inherent capabilities to discern relevant from irrelevant passages are likely already present in pre-trained BERT, and DPR-style training does not substantially enhance these discriminative features.
@@ 3 Knowledge Decentralization in DPR-Trained Models
The Riemann approximation was used due to the intractability of calculating a continuous integral. Following Dai et al. (2022), a threshold of $0.1*\max(\textrm{Attr})$ was applied to identify a coarse set of knowledge neurons [^1] Appendix A.2 demonstrates that our observations are consistent across a spectrum of thresholds.. In contrast to Dai et al. (2022), the coarse set of knowledge neurons was not refined to a fine set of knowledge neurons, as our interest is on the broader activation patterns. When the model operates, it activates both "true-positive" and "false-positive" knowledge neurons indiscriminately according to their attribution scores. The primary interest lies in how DPR training influences these activation patterns, rather than the role of specific neurons.
Answer in Top-1? # Strongly Activated Neurons Title of Top-5 Retrieval
Query Pre-trained BERT DPR-BERT Pre-trained BERT DPR-BERT Pre-trained BERT DPR BERT
where is the most distortion on a robinson projection 220 1323 Circle of latitude, Scale-invariant feature transform, Line moiré, Theil–Sen estimator, Pole splitting Robinson projection, Robinson projection, Arthur H. Robinson, Robinson projection, Arthur H. Robinson
who is the chief legal advisor to the government 65 831 Jimly Asshiddiqie, Judicial system of Iran, Comptroller General of the State Administration, Jimly Asshiddiqie, Law of Kosovo Attorney General of India, Attorney general, Attorney General of India, K. K. Venugopal, Attorney General of India
what type of government does kenya have 2018 74 287 Government of Kenya, Abundant Nigeria Renewal Party, 2007–08 Kenyan crisis, Independent Electoral and Boundaries Commission, Kingdom of Kongo Government of Kenya, Politics of Kenya, Government of Kenya, Government of Kenya, Government of Kenya
are pure metals made of atoms or ions 69 1268 Alloy, Common attributes, Metal, Resonance ionization, Alloy Properties of metals, metalloids and nonmetals, Properties of metals, metalloids and nonmetals, Solid, Metal, Metal
who is the bad guy in lord of the rings 100 533 Millennium Earl, The Sword of Shannara, Eye of Ra, The Enchanted Apples of Oz, Ys I & II Saruman, Saruman, Sauron, Morgoth, Legolas
when were manatees put on the endangered list 42 1522 Ivory trade, Namib Desert Horse, Endangered Species Act of 1973, Iriomote cat, Bile bear Manatee conservation, Endangered Species Act of 1973, Endangered Species Act of 1973, Manatee conservation, Endangered Species Act of 1973
when did wesley leave last of the summer wine 38 1024 Naif (band), Aiden, Queensrÿche, Josef Brown, Matthew Stocke Gordon Wharmby, Gordon Wharmby, Brian Wilde, Cory Monteith, Last of the Summer Wine
when did mozart compose his first piece of music 74 364 Wolfgang Amadeus Mozart, Der Messias, Life of Franz Liszt, Die Entführung aus dem Serail, Quattro versioni originali della R̈itirata notturna di Madrid¨ Wolfgang Amadeus Mozart, Wolfgang Amadeus Mozart, Leopold Mozart, Wolfgang Amadeus Mozart, Wolfgang Amadeus Mozart
Table 2: This table presents example queries alongside the corresponding model retrievals and the count of strongly activated neurons for both pre-trained and DPR-trained BERT. Notably, DPR training consistently increases the number of strongly activated neurons. Additionally, the retrievals, even when DPR does not retrieve the correct passage in the top-1 retrieval, are much more focused and targeted to the asked query. In contrast, pre-trained BERT’s retrievals are much more varied and sporadic. This is likely because in pre-trained BERT each neuron that is activated is responsible for more information and the model has no fine-grained path to follow for specific information like it does after DPR training.
@@
Figure 1 illustrates the impact of DPR training on BERT’s neuron activations, charting the attribution score of every neuron across both the intermediate and output linear layers within each transformer block for the query model [^2] Appendix A.1 shows that the observations made in this section also hold for the context model.. DPR-trained BERT has more activated neurons in the intermediate layer of each block. The output layer, on the other hand, maintains a consistent number of activations at each transformer block compared to pre-trained BERT, and in the earlier layers DPR-trained BERT activates fewer neurons in the output layers. Previous studies have conceptualized intermediate layers as "keys" and the output layer as the "value" Geva et al. (2021). This suggests that DPR training expands the set of "keys" available to access a given volume of semantic knowledge while decreasing the accessible volume of syntactic knowledge, embodying a decentralization strategy for semantic knowledge. Rather than relying on a single, highly precise key to unlock some knowledge, DPR allows for the use of multiple, somewhat less precise keys. This underscores DPR training’s primary goal: to modify the model’s method of knowledge access without altering the stored knowledge itself. These multiple pathways enable morphologically distinct but semantically related text to trigger the same knowledge or collections of facts, thus making retrieval possible.
Table 2 demonstrates the effects of DPR training through performing retrieval with various queries and the full corpus of 21M Wikipedia passages. Across all instances, DPR training increases the number of strongly activated neurons indicating the existence of more pathways in the network allowing for better access to the information needed to perform retrieval. When examining the titles of the retrieved passages, a marked difference is revealed between pre-trained BERT and DPR BERT. Pre-trained BERT’s retrievals are often disparate, aligning with the query in some instances while seemingly unrelated in others. This inconsistency indicates that successful retrievals by pre-trained BERT may hinge on the activation pattern precisely aligning with the relevant article. On the other hand, DPR-BERT, consistently retrieves passages that are topically related to the query, even if they are not the exact best match, reflecting a better ability to home in on pertinent information. By having more neurons responsible for each query the model has more fine-grained control to find relevant passages, even if it is not the most relevant passage. In the cases where it was not able to navigate to the exact correct passage it is possible that the knowledge needed to discern between the correct and incorrect passage in the article is not in the model.
@@ 4.2 Knowledge Removal
Transformer-Patch 0.16 689 0.87 183
MalMen 0.11 721 0.81 261
Mend 0.11 722 1.00 252
Table 4: This table presents the outcomes of the knowledge removal experiments. The "Probing Removed" column is the percentage of the total facts that were successfully removed from BERT. The "DPR Removed" column is the percentage of those facts that were detected after DPR training.
@@ 5 Related Works
DPR addresses the challenge of matching a query with the most relevant passages from a knowledge base Karpukhin et al. (2020). This approach employs dual encoders—one encoder for the passages and another for the query—and utilizes a distance metric, such as the inner product, to identify the passages closest to the query. Inspired by Siamese networks Bromley et al. (1993), DPR represents the first fully neural architecture to outperform the BM25 algorithm Robertson and Zaragoza (2009). Since then, there have been quite a few improvements in how to train DPR-style models. Methods like RocketQA improve DPR by employing cross-batch negatives and training the network on more difficult hard negatives Qu et al. (2021). Dragon focuses on novel data augmentation and supervision strategies Lin et al. (2023). Contriever also employs a greater number of hard-negatives and data-augmentation methods in addition to pre-training the model on the inverse cloze task Izacard et al. (2022). MVR generates multiple views for each document to allow for multiple diverse representations of each of them Zhang et al. (2022). ColBERT employs token embeddings for more fine-grained matching Khattab and Zaharia (2020). REALM leverages feedback from the reader component to jointly train the retriever with the reader Guu et al. (2020). Other methods distill knowledge from the reader to the retriever Izacard and Grave (2020); Reichman and Heck (2023). Additionally, efforts in query augmentation or generation aim to better synchronize the query with the document encoder Ma et al. (2023); Wang et al. (2023); Shao et al. (2023); Gao et al. (2023). Despite these different enhancements, each method builds upon the DPR framework discussed in this paper.
@@ 6 Conclusion
To reveal possible avenues for improving RAG systems, this paper set out to study the purpose served by DPR-style fine-tuning and how DPR-trained BERT operates. Through linear probing in Section 2, alongside experiments where we added and removed knowledge from pre-trained BERT in Section 4, we determined that BERT does not appear to acquire new information through DPR fine-tuning. Instead, we observed that the efficacy of retrieval hinges on the activation of shared facts/memories between the BERT models used to encode the query and the context passages. This mechanism implies that incorrect retrieval could occur if a query or context passage inadvertently activates irrelevant or incorrect memories in BERT. Moreover, the absence of necessary facts or webs of knowledge within the model hampers its ability to retrieve information.
@@
In the most fundamental sense, DPR achieves its namesake function—it retrieves, locating and returning relevant context to the user given a query. Yet, as our evidence suggests, DPR models appear constrained to retrieving information based on the knowledge that preexists within their parameters, either innately or through augmentation. This operational boundary delineates a significant caveat: facts must already be encoded within the model for useful context to be accessible by retrieval. Absent these facts or their associative networks, retrieval seems to falter. Thus, if retrieval is understood as the capacity to recall or recognize knowledge already familiar to the model, then indeed, DPR models fulfill this criterion. However, if we extend our definition of retrieval to also encompass the ability to navigate and elucidate concepts previously unknown or unencountered by the model—a capacity akin to how humans research and retrieve information—our findings imply that DPR models fall short of this mark.